OpenStudy (anonymous):

Order the set of numbers from least to greatest: 2.71 repeating 71, 2 and 3 over 4, square root 5, 5 over 2 2.71 repeating 71, 2 and 3 over 4, square root 5, 5 over 2. square root 5, 5 over 2, 2.71 repeating 71, 2 3 over 4 5 over 2, square root 5, 2.71 repeating 71, 2 3 over 4 2 3 over 4, 2.71 repeating 71, 5 over 2, square root 5

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

what do you mean by: 2.71 repeating 71

OpenStudy (anonymous):

try to calculate each number. Use calculator or just do it in your head. Start with 5/2=?

OpenStudy (anonymous):

\(\sqrt{5}=2.23606...\) \(\huge2\frac{3}{4}=2.75\)

OpenStudy (anonymous):

now just order this